# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 3.1 Eight classes of grand-coalition-first action models
A grand-coalition-first action model is a tuple $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
Definition 9 (Outcome functions and availability functions of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}.$
Definition 10 (Successor functions of grand-coalition-first action models).
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Define the successor function $\mathtt{suc}$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}.$
For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$. It is easy to verify the following fact:
Fact 2.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathtt{suc}(s)$.
Fact 3.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})$.
Note that neither of the converses of the two statements holds. Figure 1 indicates a counterexample for them.
Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts. A grand-coalition-first action model is indicated in the figure, where we suppose $\mathtt{AG}=\{a,b\}$. It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{a}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
@@
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})}$ is a singleton.
@@ 3.2 Eight axiomatic systems
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
@@
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$.
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
@@
Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [8]).
@@
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? All this is yet unknown.
@@ 4.1 Outcome functions are compositional in many situations
We say outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
Outcome functions in grand-coalition-first action models are not compositional. Figure 2 gives a counterexample for this.
@@
Figure 2: This figure offers a counterexample that outcome functions in grand-coalition-first action models are not generally compositional. We suppose $\mathtt{AG}=\{a,b\}$. Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted above, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$. In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$ In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$. Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$. This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
What follows are two examples, and the second is from [3].
@@
Adam and Bob control a ship lock with two doors: a front door and a back door. Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make the two doors open at the same time. This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1. Here, $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{f})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
@@
A system has two processes, $a$ and $b$. The process $a$ can assign $1$ or 00 to the variable $x$, and the process $b$ can assign $1$ or 00 to the variable $y$. When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
We want to mention that some literature, such as [6] and [7], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.

### v3

@@ 3.1 Eight classes of grand-coalition-first action models
An action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a grand-coalition-first action model if for all $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
It can be seen that for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}$ can be derived from $\mathsf{out}_{\mathtt{AG}}$, which is why we use “grand-coalition-first”.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
@@
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial if $\mathtt{av}_{\mathrm{C}}(s)\neq\emptyset$ for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic if $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})$} is a singleton for all $s\in\mathtt{ST}$ and $\sigma_{{\mathtt{AG}}}\in\mathtt{av}_{\mathtt{AG}}(s)$.
@@ 3.2 Eight axiomatic systems
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
@@
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi}$
We let the following formulas respectively correspond to $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$:
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{D}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{C}\cup\mathrm{D}]\hskip-2.0pt\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$
@@
Theorem 2 (Soundness and completeness of $\mathsf{MCL+X}$ [li_completeness_2024]).
@@
3.3 Some facts about grand-coalition-first action models
In this subsection, we present some facts about grand-coalition-first action models that will be used later.
For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$. It is easy to verify the following:
Fact 2.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
Then, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
Fact 3.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}$: $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\subseteq\mathsf{out}_{a_{1}}(s,\sigma_{{a_{1}}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{{a_{n}}})$.
Neither of the converses of the two statements holds. What follows is a counterexample for both of them:
Example 1.
Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts.
We suppose $\mathtt{AG}=\{a,b\}$. A grand-coalition-first action model is indicated in Figure 1, where $(\alpha_{i},\beta_{j})$ represents the joint action of $a$ doing $\alpha_{i}$ and $b$ doing $\beta_{j}$. It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, is equal to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{b}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, is equal to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
@@ 3.3 Some facts about grand-coalition-first action models
The class of grand-coalition-first action $\mathsf{SID}$-models is $\alpha$-representable by a class of alpha neighborhood models meeting certain properties ([pauly_modal_2002, goranko_strategic_2013]). The class of grand-coalition-first action $\mathsf{SD}$-models is $\alpha$-representable by a class of alpha neighborhood models meeting certain properties ([shi_representation_2024]). Are the other six classes of grand-coalition-first action models $\alpha$-representable in a similar sense? Are the eight classes of grand-coalition-first action models $z$-representable in a similar sense? All this is yet unknown.
@@ 4.1 Outcome functions are compositional in many situations
We say that outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
Outcome functions in grand-coalition-first action models are not compositional. What follows is a counterexample:
Example 2.
We suppose $\mathtt{AG}=\{a,b\}$. Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted in Figure 2, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$.
In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$.
In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$.
Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}$ $(s_{0},(\alpha_{1},\beta_{1}))$. This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
@@
Figure 2: This figure indicates two grand-coalition-first action models for Example 2.
There are many situations in which every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
What follows are two examples, and the second is from [alur_alternating-time_2002].
@@
Adam and Bob control a ship lock with two doors: a front door and a back door. Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make both doors open at the same time. This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 3. Here: $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed; $(\mathtt{open}\text{-}\mathtt{f},\mathtt{skip})$ represents a joint action where Adam closes the front door and Bob does nothing, and so on. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{b})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
@@
A system has two processes, $a$ and $b$. The process $a$ can assign $1$ or 0 to the variable $x$, and the process $b$ can assign $1$ or 0 to the variable $y$. When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, $b$ can either leave the value of $y$ unchanged or change it to $1$. When $y=1$, $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
Remarks
We want to mention that some literature, such as [harrenstein_boolean_2001] and [hoek_logic_2005], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.